OpenStudy (anonymous):

PLEASE HELP!!!!! Is there a hole in the equation (x^2-3x+2)/(x-2) if there is what is it?

OpenStudy (anonymous):

factor and cancel i guess

OpenStudy (anonymous):

please help me @satellite73

OpenStudy (anonymous):

if you can cancel the common factor of \(x-2\) then there will be a hole at \(x=2\) and \(y=\) whatever you get when you replace x by 3

OpenStudy (anonymous):

did you factor?

OpenStudy (anonymous):

hint, i cannot be hard, because you know one of the factors is \(x-2\) in the numerator

OpenStudy (anonymous):

when i factored i got (x-1)(x-2)

OpenStudy (anonymous):

ok good now cancel

OpenStudy (anonymous):

\[\frac{(x-1)(x-2)}{x-2}=?\]

OpenStudy (anonymous):

so i'm left with (x-1)/1 which is x-1

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

and thats all i have to do?

OpenStudy (anonymous):

hold on i am totally wrong

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

its (2,1)

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

when you cancel \(x-2\) you are getting rid of the zero in the denominator at \(x=2\) if you replace \(x\) by \(2\) you get \((2,1)\)

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

(5x^2+10x)/(-4x-8)

OpenStudy (anonymous):

and can you help me find the vertical asymptote too?

OpenStudy (anonymous):

factor a common factor of \(5x\) out of the numerator and \(-4\) out of the denominatorq

OpenStudy (anonymous):

so i get 5x(x+2)/-4(x+2)

OpenStudy (anonymous):

\[\frac{5x^2+10x}{-4x-8}=\frac{5x(x+2)}{-4(x+2)}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

leaves you with only \[-\frac{5}{4}x\]

OpenStudy (anonymous):

so i get the hole at (-2,2.5)

OpenStudy (anonymous):

no @satellite73 !!!!! why did you leave me?????

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i really need help you have NO idea

OpenStudy (anonymous):

ok yeah you got it right \((-2,2.5)\)

OpenStudy (anonymous):

how would i find the vertical asumptote?

OpenStudy (anonymous):

in those two examples there are none

OpenStudy (anonymous):

wow, how do you know?

OpenStudy (anonymous):

because you were left with no denominators when you factored and cancelled

OpenStudy (anonymous):

if you had a denominator left, the vertical asymptote would be the number that would make the denominator zero

OpenStudy (anonymous):

wasn't i left with 5x/-4? isnt -4 a denominator than?

OpenStudy (anonymous):

ok yes it is a denominator, but it is just a number

OpenStudy (anonymous):

and we need a denominator with x in it

OpenStudy (anonymous):

yes of course \[y=-\frac{5}{4}x\] is just a line

OpenStudy (anonymous):

if you had \[y=\frac{x}{x-1}\] then you would have a vertical asymptote at \(x=1\)

OpenStudy (anonymous):

one last i found the x and y intercepts to be (0,0) and (0,0) is that correct?

OpenStudy (anonymous):

i think i understand the vertical asymptote now

OpenStudy (anonymous):

for \(y=-\frac{5}{4}x\) yes

OpenStudy (anonymous):

it is a line through the origin with slope \(-\frac{5}{4}\)